# Bones of Contention: Exploring Query-Efficient Attacks Against Skeleton Recognition Systems

### v1

@@ V-G Extension to Targeted Attacks
Targeted Attacks for ISAAC-K. We also conduct targeted attacks for ISAAC-K. According to our experiments, the original search method of RayS works well enough for untargeted attacks, and its advantages in the $\ell_{\infty}$ setting are well adapted to our requirements for motion naturalness. However, RayS is known to perform poorly in targeted attacks since there is no appropriate initial search direction for the original motion. Therefore, we propose a search direction optimization here. Following gradient-based targeted attacks [56], we start the targeted attack from a motion selected in the target class and take the direction from the target motion to the original motion as the initial direction. This direction enables the search to reach the target area faster. Also, as an improvement of RayS, we do not start the search with a radius of $\infty$. Instead, we replace it with the decision boundary distance of the initial direction. This will reduce the search space to avoid redundant searches in the early stage. Overall, we start the attack from a target class motion $\mathbf{x}_{t}$, and optimize the search direction by setting $\mathbf{d}_{best}\leftarrow M\odot{\mathop{\rm sgn}}\left(\mathbf{x}_{t}-\mathbf{x}\right)$, ${r_{best}}\leftarrow{g_{res}}\left({{{\mathbf{x}}_{t}}-{\mathbf{x}}}\right)$.
@@
TABLE VIII: Attack performance comparison for targeted attacks.
@@
MS-AAGCN BASAR [20] 0.09 123.02 0.12 2.17 0.08 168.93 0.11 2.54
@@
Targeted Attacks for ISAAC-N. We only conduct untargeted attacks for ISAAC-N in this paper since devising a comprehensive attack from the perspective of non-semantic regions is not the focus of this paper. Nonetheless, we provide a somewhat feasible idea to conduct targeted attacks. Specifically, we can simply divide the human body into six regions according to joints, i.e., the left hand, the right hand, the left leg, the right leg, the spine and the head. According to the number of $N_{n}\left(=N-N_{k}\right)$ joints belonging to the above six regions, we can obtain one or two non-semantic regions and replace them as a whole, i.e., replacing the sitting posture in both leg areas with kneeling. For targeted attacks, one possible solution is to first collect sets of key joints in different regions from the target class, and then select a motion $\mathbf{x}_{t}$ from the set according to the region to be replaced by the sample to be attacked and then replace it in the following way.
${\mathbf{x}_{rep}}=\Psi\left({\mathbf{x}\odot\left({1-{M_{r}}}\right),{\mathbf{x}_{t}}\odot{M_{r}}}\right),$ (12)
where ${\mathbf{x}_{rep}}$ denotes the replaced motion, and $\Psi$ denotes the splicing operation of two motions, including the translation of the root joint and the normalization of the motion. Then, minor perturbations can be added until the motion is misclassified to the target class.
@@ VI Discussion
Long Videos as Input. We use the count-based key joint selection since videos in our dataset mostly contain only one action. If the input video is long and contains more actions, the key joints of different frames may appear different. As a possible alternative solution, the motion $\mathbf{x}$ can be split into $K$ sub-action segments $\left\{{{S_{1}},{S_{2}},...,{S_{K}}}\right\}$, where ${S_{i}}=\left\{{{\mathbf{x}_{{t_{i}}}},{\mathbf{x}_{{t_{i}}+1}},...,{\mathbf{x}_{{t_{i}}+{T_{i}}-1}}}\right\}$, $\mathbf{x}_{t_{i}}$ represents the ${t_{i}}$-th frame of the motion, ${t_{i}}$ and ${T_{i}}$ denote the first frame index and the frame number of the $i$-th segment. Then the key joints are extracted respectively among different segments but remain the same in each segment.
Other Scenarios of Inputs. For video inputs, there exist mature methods [57] for extracting keypoints from videos. In existing skeleton datasets, the 3D coordinates are estimated from the 2D data. Although keypoint extraction is becoming increasingly refined, this does not mean that more keypoints lead to more effective action recognition; rather, it poses greater challenges [58].
Different Surrogate Models. While Grad-CAM may yield varying key joints across different surrogate models, this variability is unlikely to substantially affect attack performance. This conclusion is supported by a pilot experiment that assesses the cosine similarity of the obtained key joints for eight models on NTU60(3D). As shown in Figure 13, the key joints are similar among models, indicating that changing the surrogate model has little impact on the output results. It is observed that ST-GCN achieves the highest overall key joint similarity with other models, which also provides support for choosing ST-GCN as the surrogate model in our experiments.
Poisoning Attacks. Due to the adoption of online learning in some outsourcing services, which continuously updates the model in the cloud service based on new input data, the model becomes malicious once the data are contaminated, which may cause a loss of utility of the whole pipeline.
@@ VII Conclusion
Social Impact. As an adversarial attack, our proposed method may be maliciously or unintendedly used to threaten the security of skeleton recognition systems in the real world. We hope that our work can arouse the related community to attach greater significance to the vulnerability of skeletal models, since their accuracy, as well as robustness, are crucial to widely used scenarios, such as aged care, video surveillance, etc. We believe that our discovery of a more efficient and effective sparse attack, as well as the non-semantic attack, will be accorded due significance and treated as a novel respect to help improve the robustness of skeletal models in steeply rising safety-critical systems.
@@ Acknowledgments
The authors would like to thank the anonymous reviewers for their valuable feedback that helped improve the paper. This work was supported in part by Ping An Technology and National University of Singapore.

### v2

@@ V-G Extension to Targeted Attacks
Targeted Attacks for ISAAC-K. We also conduct targeted attacks for ISAAC-K. According to our experiments, the original search method of RayS works well enough for untargeted attacks, and its advantages in the $\ell_{\infty}$ setting are well adapted to our requirements for motion naturalness. However, RayS is known to perform poorly in targeted attacks since there is no appropriate initial search direction for the original motion. Therefore, we propose a search direction optimization here. Following query-based targeted attacks [ilyas2018black], we start the targeted attack from a motion selected in the target class and take the direction from the target motion to the original motion as the initial direction. This direction enables the search to reach the target area faster. Also, as an improvement of RayS, we do not start the search with a radius of $\infty$. Instead, we replace it with the decision boundary distance of the initial direction. This will reduce the search space to avoid redundant searches in the early stage. Overall, we start the attack from a target class motion $\mathbf{x}_{t}$, and optimize the search direction by setting $\mathbf{d}_{best}\leftarrow M\odot{\mathop{\rm sgn}}\left(\mathbf{x}_{t}-\mathbf{x}\right)$ and ${r_{best}}\leftarrow{g_{res}}\left({{{\mathbf{x}}_{t}}-{\mathbf{x}}}\right)$.
@@
TABLE VIII: Attack performance comparison of targeted ISAAC-K.
@@
MS-AAGCN BASAR [diao2021basar] 0.09 123.02 0.12 2.17 0.08 168.93 0.11 2.54
@@
Targeted Attacks for ISAAC-N. We only conduct untargeted attacks for ISAAC-N in the main experiment since devising a comprehensive attack from the perspective of non-semantic regions is not the focus of this paper. Nonetheless, we provide a somewhat feasible idea to conduct targeted attacks. Specifically, we can simply divide the human body into six regions according to joints, i.e., the left hand, the right hand, the left leg, the right leg, the spine and the head. According to the number of $N_{n}\left(=N-N_{k}\right)$ joints belonging to the above six regions, we can obtain one or two non-semantic regions and replace them as a whole, i.e., replacing the sitting posture in both leg areas with kneeling. For targeted attacks, one possible solution is to first collect sets of key joints in different regions from the target class, and then select a motion $\mathbf{x}_{t}$ from the set according to the region to be replaced by the sample to be attacked and then replace it in the following way:
$\small{\mathbf{x}_{rep}}=\Psi\left({\mathbf{x}\odot\left({1-{M_{r}}}\right),{\mathbf{x}_{t}}\odot{M_{r}}}\right),$ (15)
where $\mathbf{x}_{rep}$} denotes the replaced motion, and $\Psi$ denotes the splicing operation of two motions, including the translation of the root joint and the normalization of the motion. Then, minor perturbations can be added until the motion is misclassified to the target class. We conduct targeted attacks on NTU120(3D) and choose “sit” as the target. Results in Table IX indicate that ISAAC-N can achieve certain targeted performance, thereby enhancing its applicability to more realistic threat scenarios.
TABLE IX: Attack performance for targeted ISAAC-N.
Model ISAAC-NR ISAAC-NRL ISAAC-NRA
[…]
MS-AAGCN 0.05 0.00 0.08 6.38 0.49 15.10
CTR-GCN 0.05 0.00 0.09 7.33 0.38 15.55
DG-STGCN 0.00 0.00 0.00 0.00 0.07 15.71
MS-G3D 0.08 0.00 0.13 6.38 0.45 14.82
ST-GCN 0.14 0.00 0.19 4.84 0.37 14.81
ST-GCN++ 0.13 0.00 0.17 4.41 0.45 14.89
V-H Surrogate Model Analyses
Dependency on Surrogate Model. In our pipeline, the surrogate model is used only to extract key joints that guide the placement of sparse perturbations. This step does not require the surrogate to resemble the victim model, since it merely approximates motion-sensitive regions. In realistic black-box scenarios, attackers can freely choose any open-source skeletal model for this purpose, since only the victim model is inaccessible. While Grad-CAM may yield varying key joints across different surrogate models, this variability is unlikely to substantially affect attack performance. This conclusion is supported by a pilot experiment that assesses the cosine similarity of the obtained key joints for eight models on NTU60(3D). As shown in Fig. 13, the key joints are similar among models, indicating that changing the surrogate model has little impact on the output results. It is observed that ST-GCN achieves the highest overall key joint similarity with other models, which also provides support for choosing ST-GCN as the surrogate model in our experiments.
Role of the Surrogate Model. The role of the surrogate in our method differs fundamentally from that in traditional gray-box (transfer-based) attacks. Prior work typically optimizes perturbations in a white-box manner on a surrogate and then transfers them to the victim model, which makes the attack highly dependent on the surrogate architecture and often results in low transferability, e.g., perturbations optimized for the surrogate may fall back inside the decision boundary of the victim model and fail to cause misclassification. In contrast, our method treats the surrogate model solely as a tool for extracting key joints used to guide the perturbation process. The actual attack optimization is carried out entirely in the black-box setting by directly querying the victim model. This results in a pure black-box attack that does not rely on architectural similarity between the surrogate and the victim model, and is thus significantly more robust to model mismatch.
@@ VI Discussion
Input Feature Dimension. In skeleton-based action recognition, most mainstream models adopt a consistent input format, typically 25 joints, due to the standardization established by the NTU 60/120 datasets. Captured using Kinect V2, which detects 25 body joints, these datasets have become the dominant benchmark, and most models, including GCN-based and ViT-based architectures, are designed accordingly. Therefore, the issue of dimensional mismatch between surrogate and victim models rarely arises. That said, if, in the future, new sensing hardware springs out and leads to dominant datasets with fine-grained skeletal structures (e.g., 30 joints), we believe that new standardized formats are likely to emerge, with models adjusted to support the updated input. Even in the absence of a unified format, a surrogate model can be adapted to match the victim model’s input. For example, one could start with a well-performing 25-joint model, adjust its early layers to accept 30-joint data, and train it to mimic the outputs of the victim model. In this way, the surrogate model remains aligned in structure and output behavior, allowing key joint extraction to proceed as in our proposed attack.
Long Videos as Input. We use the count-based key joint selection since videos in our dataset mostly contain only one action. If the input video is long and contains more actions, the key joints of different frames may appear different. As a possible alternative solution, the motion $\mathbf{x}$ can be split into $K$ sub-action segments $\left\{{{S_{1}},{S_{2}},...,{S_{K}}}\right\}$, where ${S_{i}}=\left\{{{\mathbf{x}_{{t_{i}}}},{\mathbf{x}_{{t_{i}}+1}},...,{\mathbf{x}_{{t_{i}}+{T_{i}}-1}}}\right\}$, $\mathbf{x}_{t_{i}}$ represents the $t_{i}$}-th frame of the motion, $t_{i}$} and $T_{i}$} denote the first frame index and the frame number of the $i$-th segment, respectively. Then the key joints are extracted respectively among different segments but remain the same in each segment.
Other Scenarios of Inputs. For video inputs, there exist mature methods [zheng2023deep] for extracting keypoints from videos. In existing skeleton datasets, the 3D coordinates are estimated from the 2D data. Although keypoint extraction is becoming increasingly refined, this does not mean that more keypoints lead to more effective action recognition; rather, it poses greater challenges [song2021human], and can serve as a potential direction for future research.
Social Impact. As an adversarial attack, our proposed method may be maliciously or unintentionally used to threaten the security of skeleton recognition systems in the real world. We hope that our work can arouse the related community to attach greater significance to the vulnerability of skeletal models, since their accuracy, as well as robustness, are crucial to widely used scenarios, such as aged care, video surveillance, etc. We believe that our discovery of a more efficient and effective sparse attack, as well as the non-semantic attack, will be accorded due significance and treated as a novel respect to help improve the robustness of skeletal models in steeply rising safety-critical systems.
@@ Acknowledgments
The authors thank the anonymous reviewers for their valuable feedback that helped improve the paper. This work was supported in part by National University of Singapore (No.A-8003872-00-00), the University of Hong Kong and Ping An Technology. This work was also partially supported by the NSFC for Young Scientists of China (No.62202400) and the RGC for Early Career Scheme (No.27210024). Any opinions, findings, or conclusions expressed in this material are those of the authors and do not necessarily reflect the views of NSFC and RGC.